Operations Report – March 23rd

Crew 207 Operations Report 23 Mar 2019

SOL: 7

Name of person filing report: Tom Baldwin

Non-nominal systems: Generator: Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.

Generator:

Generator (hours run): (TBD, attempted contact with Dave.) hours

Generator was turned off at (TBD, attempted contact with Dave), SOC (TBD) %

Diesel Reading – 50 %

Propane Reading – 62 %

Ethanol Free Gasoline – 2 gallons.

Water (auxiliary tank) – Not in use

Water (static tank) – 160 gallons

Auxiliary to Static tank transfer– Not applicable

Gallons transferred: Not Applicable

Water in Green Hab – 148 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: 142050.1 gallons

Toilet tank emptied: No

Deimos rover used: No, not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Not used

Hours: 85.1 hours

Beginning charge: 100 %

Ending charge: 100 %

Currently charging: Yes

Opportunity rover used: Not used

Hours:54.9hours

Beginning charge: 100 %

Ending charge: 95 %

Currently charging: Yes (now 100%)

Curiosity rover used: Used

Hours: 84.7 hours

Beginning charge: 100 %

Ending charge: 97 %

Currently charging: Yes (now 100%)

Notes on rovers: not used

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): N/A

Reason for use: Not used

Oil Added?: N/A

ATV Fuel Used: N7A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar: None.

CrewCar used and why, where? David reports that Crew Car is back in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: EVA suits normal, with all batteries greater than 13 volts, (including the three used for EVAs today).

Summary of Hab operations: Checked SOC of the electrical system.

08:00 79%, generator on, fault light on. Reported to David.
12:04 92%, generator off, fault light off.
16:00 74% generator off, fault light off. Reported to David.

Summary of Green Hab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Questions, concerns and requests to Mission Support: None